Tech Staff Meeting Notes April 7, 2005

Current Top Priorities are:GPS, Station Maintenance, I-40 bridge

There are four levels of priority.
PrioritiesWish to try to get these done before the next meeting (yeah right)
Second TierFit this in too if an opportunity arises. May be good to do instead of priorities but check with Mitch. Likely to be a priority in a previous and/or subsequent month
As Time PermitsFit these in when you can but not instead of first or second priorities
OtherStuff we hope to get to and don't want to forget about


Reelfoot. On hold until the state decides to finish the remodel. Given this uncertain it is desireable to do something longer-term temporary (how's that for a wishy washy phrase) to avoid a large monitoring gap. Install some air-cells and do routine visits to replace them and download data.
Mineral Area College. Need to replace it with an IP ready receiver. The snag here is that all the receivers need to get serviced under an RMA and Smalley doesn't want to install any until that is done. DONE.
Bob will acquire 4 raydomes to replace damaged ones.
Chris will integrate routine maintenance trip to all GPS sites for visual check, kudzu control, battery check, repaint masts, etc. Also check them out for best way to build/install a cover to help control temperature.
Steele and Charleston are still dialup. What's the best way to get them real-time? Charleston might be line of site to NMAD. At 1sps they generate about 32M/day.

ANSS. E. St Louis, IL. done.
Evansville, IN. next week.
Began swapping all sensors; Filipcic is taking the lead and keeping records. First 3 went to Wurth and are back at DTA. Next three will go to Parker. Filipcic will try to get a time frame and if its too long, Parker will get Charleston going again without swapping.

Routine Maintenance. Mitch needs to make sure he leaves few weeks this summer to hit all stations.

I-40 Bridge. Power failure fixed. AC will get switched to the new system in mid-April.
One digitizer is not networking; probably the network card.
Remote reset will be improved hopefully next along with getting HDBT real-time.
Many loose ends to complete when we can get a lane closure; probably about 2 or 3 weeks of work remaining.
Need a development/test DAS for Brewer.
Steiner provided the DSP filter description to Mitch; might be able to solve for poles and zeros.
Sharam is preparing a proposal to put borehole sensors on Mud Island near HDBT.

RSCM was destroyed by vandals. This is the third time. It is a quiet and optimally located within the network. Need to replace and/or move if a suitable alternative can be found within a few kilometers.

Second Tier Priorities

SWET. Backonline but hardware needs replacing. Withers contacted Benz and upgrades are going to happen but we don't know when (probably not soon).

TZTN. Done. Mitch will contact Alena regarding episensor so we can get SLU's returned to them (no word yet). Parker will install the gate at his leisure.

BB network upgrade. CTS/RTS are not enabled on the dm24 which required a redesign of the Qualtrons. They have a new idea and will give it two more months (after which they cut their losses and bail). Should work though.
Will do a huddle test in the bb test vault.
Dave at Utah has freewave demuxing software that could simplify things at the nodes.

FDL. Not discussed at this meeting.

CVTN and ASAR optical isolation. Need to upgrade to year 3 hardware. Hardware ready. Filipcic needs to install.

Blohole tree fell down. Reckon we really do need to reinstall now.

John has 3 stations in east TN and 2 in North Carolina to install. Would like to get some help over there to do this and to get LRNC back online.

McGoldrick will acquire station decals to provide name and contact information.

As Time Permits

Battery charging station. Need a hood and eyewash.

Cowcamps pit is done. A tower is being constructed at the top of the mountain that have handshake permission to get on.

Pressure wash and paint MKTA equipment hut.


Greg provided a todo list that is integrated into this list but for reference, the original is here.

FDL flooding. Done.

MPHB. Withers needs to determine current orientation so we can pull it and orient correctly. Helicorders look strange. Need to verify the data.

BB Test Facility. Need to finish wiring and electronics (see note above on BB upgrade huddle test).

Timing. Steiner has an easy way to generate a datastream that goes high when lock is lost. Need a mask to decide what streams to use. Need mechanism to produce alarms. BB stations more difficult. How do we automate the status stream or even get it back to Memphis for logging? (Save it to file with Scream?). Note, new version of firmware sends status to both flash and stream. Hence, information will be in gcf2ew logs, but how do we monitor that? Maybe not a huge deal with broadbands since both Bollwerk and Brewer regularly monitor. Would still be good to have a record though.

North Carolina station refurbish. Keep node at LRNC. Modify to be the same as a New Madrid node and minimize internet bandwidth required.

VBMS. Proposal is under review and on hold pending vault decision.

GAI. BB-13 from Gainesville will be part of the huddle test.

Langston student projects

HD-8. GA Tech crew will be here in about a month to do more testing. Langston/Bodin may also want to do a seismic line. Langston/Bodin/Withers are providing coordination. Standard rules for visiting stations will be applied. We should have a list and schedule soon.

McNamara PDF. Filipcic acquired the software and is reviewing the documentation.